For each of the reactions, calculate the mass (in grams) of the product formed when 15.01 g of the underlined reactant completely reacts. Assume that there is more than enough of the other reactant. 2K(s)+Cl2(g)−−−−−→2KCl(s) 2K(s)+Br2(l)−−−−−→2KBr(s) 4Cr(s)+3O2(g)−−−−−→2Cr2O3(s) 2Sr(s)−−−−+O2(g)→2SrO(s)
Answer #1

1)

Molar mass of Cl2 = 70.9 g/mol

mass of Cl2 = 15.01 g

mol of Cl2 = (mass)/(molar mass)

= 15.01/70.9

= 0.2117 mol

According to balanced equation

mol of KCl formed = (2/1)* moles of Cl2

= (2/1)*0.2117

= 0.4234 mol

Molar mass of KCl,

MM = 1*MM(K) + 1*MM(Cl)

= 1*39.1 + 1*35.45

= 74.55 g/mol

mass of KCl = number of mol * molar mass

= 0.4234*74.55

= 31.57 g

Answer: 31.57 g

2)

Molar mass of Br2 = 159.8 g/mol

mass of Br2 = 15.01 g

mol of Br2 = (mass)/(molar mass)

= 15.01/1.598*10^2

= 9.393*10^-2 mol

According to balanced equation

mol of KBr formed = (2/1)* moles of Br2

= (2/1)*9.393*10^-2

= 0.1879 mol

Molar mass of KBr,

MM = 1*MM(K) + 1*MM(Br)

= 1*39.1 + 1*79.9

= 119 g/mol

mass of KBr = number of mol * molar mass

= 0.1879*1.19*10^2

= 22.36 g

Answer: 22.36 g

3)

Molar mass of O2 = 32 g/mol

mass of O2 = 15.01 g

mol of O2 = (mass)/(molar mass)

= 15.01/32

= 0.4691 mol

According to balanced equation

mol of Cr2O3 formed = (2/3)* moles of O2

= (2/3)*0.4691

= 0.3127 mol

Molar mass of Cr2O3,

MM = 2*MM(Cr) + 3*MM(O)

= 2*52.0 + 3*16.0

= 152 g/mol

mass of Cr2O3 = number of mol * molar mass

= 0.3127*1.52*10^2

= 47.53 g

Answer: 47.53 g

D)

Molar mass of Sr = 87.62 g/mol

mass of Sr = 15.01 g

mol of Sr = (mass)/(molar mass)

= 15.01/87.62

= 0.1713 mol

According to balanced equation

mol of SrO formed = moles of Sr

= 0.1713 mol

Molar mass of SrO,

MM = 1*MM(Sr) + 1*MM(O)

= 1*87.62 + 1*16.0

= 103.62 g/mol

mass of SrO = number of mol * molar mass

= 0.1713*1.036*10^2

= 17.75 g

Answer: 17.75 g